OpenStudy (anonymous):

give a combinatorial argument to prove the identity k(n,k)=n(n-1,k-1) I am looking at pascals identity right now to try and prove, but I am unsure how to actually work it except to explain the theorem ..... :(

OpenStudy (anonymous):

R.H.S=n(n-1,k-1)

OpenStudy (anonymous):

\[=n \left(\begin{matrix}n-1 \\k-1\end{matrix}\right)=n \frac{ \left( n-1 \right)! }{ \left\{ \left( n-1 \right)-\left( k-1 \right) \right\}!\left( n-1\right)!}\]

OpenStudy (anonymous):

\[= \frac{n*\left( n-1 \right)! }{\left( n-k \right)!\left( k-1 \right)! }*\frac{ k }{k }\]

OpenStudy (anonymous):

\[=\frac{ k \left( n! \right) }{\left( n-k \right)!*k! }=k \left(\begin{matrix}n \\ k\end{matrix}\right)\] =L.H.S

OpenStudy (anonymous):

THANK YOU!! i didn't even think to break it into its factorial.... i totally should have... but that helped ALOT!!!

OpenStudy (anonymous):

YW